OpenStudy (anonymous):

if the measure of one of the two complementary angles is 3 more then twice the measure of the other find the measure of each angle. fans 222222

OpenStudy (anonymous):



Directrix (directrix):

Complementary angles sum to 90. Let x be one of the two angles. Then, 90 - x is the measure of the second angle.

Directrix (directrix):

one angle is 3 more then twice the measure of the other x = 2*(90 - x) +3 Solve this equation for x and post what you get, okay?

OpenStudy (anonymous):

180 - 2x + 3?

OpenStudy (anonymous):

but whats x

Directrix (directrix):

You expanded the right side of the equation and ignored the left side. x = 2*(90 - x) +3 x = 180 - 2x + 3 Add 2x to both sides and post what you get.

OpenStudy (anonymous):